### v1

@@ 3 Estimation
We consider a two-step estimation procedure to derive the loading vectors and latent factors. This approach begins with initialization through randomized composite PCA, followed by an iterative refinement step utilizing an iterative simultaneous orthogonalization procedure.
@@
where $\Theta_{ij}={\mathbb{E}}\left[f_{it}f_{jt}\right]$. Its sample analogue, denoted as $\widehat{\Sigma}$, is computed as the average outer product over $T$ observations:
@@
where $\Theta=\mathbb{E}F_{t}F_{t}^{\top}$, $F_{t}=(f_{1t},\cdots,f_{rt})^{\top}$. This unfolding enables classical PCA estimation if the columns of the loading matrix $A$ are orthogonal. Our framework accommodates general non-orthogonal $a_{i}$’s and hence the PCA procedure introduces a bias component, which motivates the second stage refinement. The accuracy of the PCA estimator hinges on the maximum correlation among the loading vectors. When the additional orthogonality condition is imposed as in Babii et al., (2023), the maximum correlation reduces to 00 and hence bias disappears. The first step, termed intializaiton via randomized composite PCA, is detailed in Algorithm 1.
To further relax the eigengap assumption imposed in Babii et al., (2023) and Han et al., (2023), we incorporate randomized projection into our composite PCA approach (Procedure 2). Random projection, also known as random slicing (Anandkumar et al., 2014b, ; Sun et al.,, 2017) is a well-recognized initialization method in noiseless tensor CP decomposition, which accommodates repeated eigenvalues. We extend this approach to the tensor CP factor model.
Input : The observations ${\cal Y}_{t}\in\mathbb{R}^{d_{1}\times\cdots\times d_{K}}$, $t=1,...,T$, the number of factors $r$, small constant $10<c_{0}<$ 1.
@@
2Evaluate $\widehat{\Sigma}$ in (5), and unfold it to $d\times d$ matrix $\widetilde{\Sigma}$.
@@
Input : Noisy tensor $\Xi\in\mathbb{R}^{d_{1}\times\cdots\times d_{K}\times d_{1}\times\cdots\times d_{K}}$, rank $s$, number of random projections $L$, tuning parameter $\nu$.
@@
2for $\ell=1$ to $L$ do
3       Randomly draw a $d_{1}\times d_{1}$ Gaussian matrix $\theta$ whose entries are i.i.d. $N(0,1)$.
4      Compute $\Xi\times_{1}\times_{K+1}\theta$ and compute its leading singular value and left singular vector $\eta_{\ell},\widetilde{u}_{\ell}$.
5      Compute $\widetilde{a}_{\ell k}$ as the top left singular vector of ${\rm mat}_{k}(\widetilde{u}_{\ell})\in\mathbb{R}^{d_{k}\times(d/(d_{k}d_{1}))}$, for all $2\leq k\leq K$.
6      Compute $\widetilde{a}_{\ell 1}$ as the top left singular vector of $\Xi\times_{k=2}^{K}\widetilde{a}_{\ell k}\times_{k=K+2}^{2K}\widetilde{a}_{\ell,k-K}$.
@@
Following initialization, we refine the estimation using an iterative simultaneous orthogonalization procedure (Algorithm 3). This step aims to enhance estimation accuracy and extract latent factors. The procedure is motivated by the vector factor structure of the denoised ${\cal Y}_{t}$:
${\cal Z}_{t,ik}=f_{it}a_{ik}+{\cal V}_{t,ik},$ (8)
@@
$B_{k}=A_{k}(A_{k}^{\top}A_{k})^{-1}=(b_{1k},...,b_{rk})\in\mathbb{R}^{d_{k}\times r}$, $A_{k}=(a_{1k},\ldots,a_{rk})\in\mathbb{R}^{d_{k}\times r}$, and we have used the fact that $b_{ik}$ is orthogonal to all $a_{jk},j\neq i$ by construction. Note that the orthogonalization projection, which takes place in all except the $k$th mode simultaneously in each computational iteration, transforms the tensor ${\cal Y}_{t}$ to a $d_{k}\times 1$ vector, reducing dimensions and noise substantially. This transformation enables easy and accurate estimation of the classical vector factor model in equation (8).
In practice, we don’t observe $b_{ik}$ and iterations can be applied to update the estimations. Given the previous estimates $\widehat{a}_{ik}^{(m-1)}$, where $m$ is the iteration number, ${\cal Y}_{t}$ can be denoised via
@@
for $t=1,...,T$, and consequently, updated loading vectors $\widehat{a}_{ik}^{(m)}$ are obtained through eigenanalysis based on the contemporary covariance $\widehat{\Sigma}({\cal Z}_{1:T,ik}^{(m)})=\frac{1}{T}\sum_{t=1}^{T}{\cal Z}_{t,ik}^{(m)}{\cal Z}_{t,ik}^{(m)\top}$. The iteration continues until convergence or the maximum number of iterations is reached.
Input : The observations ${\cal Y}_{t}\in\mathbb{R}^{d_{1}\times\cdots\times d_{K}}$, $t=1,...,T$, the number of factors $r$, the warm-start initial estimates $\widehat{a}_{ik}^{(0)}$, $1\leq i\leq r$ and $1\leq k\leq K$, the tolerance parameter $\epsilon>0$, and the maximum number of iterations $M$.
@@ 10                  Given previous estimates $\widehat{a}_{ik}^{(m-1)}$, calculate
for $t=1,...,T$. Let $\widehat{\Sigma}\left({\cal Z}_{1:T,ik}^{(m)}\right)=\frac{1}{T}\sum_{t=1}^{T}{\cal Z}_{t,ik}^{(m)}{\cal Z}_{t,ik}^{(m)\top}$.
@@
$\widehat{f}_{it}$ $={\cal Y}_{t}\times_{k=1}^{K}\widehat{b}_{ik}^{(m)\top},\quad i=1,...,r,\ \ t=1,...,T,$
@@
Algorithm 3 Iterative Simultaneous Orthogonalization (ISO)
@@
Input : The observations ${\cal Y}_{t}\in\mathbb{R}^{d_{1}\times\cdots\times d_{K}}$, $t=1,...,T$, the upper bound of the number of factors $r_{\max}$.
@@
2Evaluate $\widehat{\Sigma}$ in (5), and unfold it to $d\times d$ matrix $\widetilde{\Sigma}$, i.e. $\widetilde{\Sigma}=\operatorname{\hbox{\rm mat}}_{[K]}(\widehat{\Sigma})$.
@@
Input : The observations ${\cal Y}_{t}\in\mathbb{R}^{d_{1}\times\cdots\times d_{K}}$, $t=1,...,T$, the upper bound of the number of factors $r_{\max}$.
@@ 5      Obtain $\widehat{r}_{k}$ by
Another closely related approach is tensor PCA proposed in Babii et al., (2023). They consider a CP tensor factor model with orthogonal loading vectors. Unlike Tucker factor models, the identification of CP factor models does not necessarily require orthogonality. Applying tensor PCA to models with non-orthogonal loadings introduces a bias component of higher order than our first-stage randomized composite PCA. Even when the loadings are orthogonal, our contemporary variance-based iterative estimation exhibits a faster convergence rate than tensor PCA due to dimension and noise reduction. A comparison of our estimator with the autocovariance-based estimator and tensor PCA through simulation will be presented in Section 5.
@@ 4 Theory
In this section, we delve into the statistical attributes of the algorithms introduced previously. Our theoretical framework offers guarantees for consistency and outlines the statistical error rates for estimating the factor loading vectors $a_{ik}$, where $1\leq i\leq r,1\leq k\leq K$, given certain regularity conditions. Considering that the loading vector $a_{ik}$ can only be identified with a change in sign, we employ
$\|\widehat{a}_{ik}\widehat{a}_{ik}^{\top}-a_{ik}a_{ik}^{\top}\|_{2}=\sqrt{1-(\widehat{a}_{ik}^{\top}a_{ik})^{2}}=\sup_{\boldsymbol{z}\perp a_{ik}}|\boldsymbol{z}^{\top}\widehat{a}_{ik}|$
to quantify the discrepancy between $\widehat{a}_{ik}$ and $a_{ik}$.
@@
Recall $F_{t}=(f_{1t},...,f_{rt})^{\top}$, $\Theta=\mathbb{E}(F_{t}F_{t}^{\top})$ and $\lambda_{i}=\lambda_{i}(\Theta)$ for $1\leq i\leq r$. Assume $\lambda_{1}\geq\lambda_{2}\geq\cdots\geq\lambda_{r}>0$. For any $v\in\mathbb{R}^{r}$ with $\|v\|_{2}=1$,
$\max_{t}\mathbb{P}\left(\left|v^{\top}\Theta^{-1/2}F_{t}\right|\geq x\right)\leq c_{1}\exp\left(-c_{2}x^{\gamma_{1}}\right),$ (14)
@@
Assume the factor process $f_{it},1\leq i\leq r$, is stationary and strong $\alpha$-mixing in $t$. The mixing coefficient satisfies
@@
Assumption 4.1 aligns closely with the noise conditions presented in seminal works such as Bai and Ng, (2002), Bai, (2003), Lam et al., (2011), Lam and Yao, (2012), and others within the factor model literature. For simplicity, we assume that the noise tensor remains independent across time $t$, allowing for weak cross-sectional dependence. While incorporating weak temporal correlation among the noise, as suggested by Bai and Ng, (2002), is plausible, it substantially complicates our theoretical analysis. Therefore, we defer this exploration to future research. Nonetheless, our simulation studies demonstrate the robust performance of the proposed methods even under conditions of weak temporal dependence.
Assumption 4.2 ensures the unique identification of all factor loading vectors $a_{ik}$ up to sign changes. Unlike the eigen decomposition of a matrix, if some $\lambda_{i}$ are equal, the estimation of the loading vectors $a_{ik}$ isn’t subject to rotational ambiguity but only to the signed permutation of loading vectors. Furthermore, Assumption 4.2 specifies that the tail probability of $f_{it}$ must exhibit exponential decay. Specifically, when $\gamma_{1}=2$, it implies that $f_{it}$ follows a sub-Gaussian distribution.
@@
Recall $A_{k}$ defined in equation (10) with $a_{ik}$ as its columns, and $A_{k}^{\top}A_{k}=(\sigma_{ij,k})_{r\times r}$. As $\sigma_{ii,k}=\|a_{ik}\|_{2}^{2}=1$, the correlation among columns of $A_{k}$ can be measured by
@@
to measure the correlation of the matrix $A=(a_{1},\ldots,a_{r})\in\mathbb{R}^{d\times r}$ with $a_{i}=\mbox{vec}(\otimes_{k=1}^{K}a_{ik})$ and $d=\prod_{k=1}^{K}d_{k}$. Let $\delta_{\max}=\max\{\delta_{1},\cdots\delta_{K}\}$.
@@
Suppose Assumptions 4.1, 4.2, 4.3 hold. Let $1/\gamma=2/\gamma_{1}+1/\gamma_{2}$, and $\delta<1$ with $\delta$ defined in (17). Assume $T\leq C\exp\left(d^{\vartheta/(2\vartheta+4)}\right)$ and $T\leq C\exp\left((d/r)^{\gamma_{1}/2}\right)$.
(i). The eigengaps satisfy $\min\{\lambda_{i}-\lambda_{i+1},\lambda_{i}-\lambda_{i-1}\}\leq c\lambda_{r}$ for all $1\leq i\leq r$, with $\lambda_{0}=\infty,\lambda_{r+1}=0$, and $c$ is sufficiently small constant. With probability at least $1-T^{-C_{1}}-d^{-C_{1}}$, the following error bound holds for the estimation of the loading vectors $a_{ik}$ using Algorithm 1,
@@
$\phi^{(0)}$ $=\lambda_{1}\left(\sqrt{\frac{r+\log T}{T}}+\frac{(r+\log T)^{1/\gamma}}{T}\right)+\sqrt{\frac{\lambda_{1}d\log d}{T}}+\frac{\sqrt{\lambda_{1}}(d\log d)}{T}+1.$ (19)
@@
The first term of the upper limit in (18) and (20) arises due to the loading vectors $a_{ik}$ not being orthogonal, which may be seen as bias. Meanwhile, the subsequent term in (18) is derived from a concentration bound concerning random noise, thereby being describable as a form of stochastic error. The condition $T\leq C\exp\left(d^{\vartheta/(2\vartheta+4)}\right)$ and $T\leq C\exp\left((d/r)^{\gamma_{1}/2}\right)$ is imposed primarily for convenience. Eliminating this condition would result in a more complex convergence rate.
When the eigengap condition is not met, we employ randomized projection to determine the statistical convergence rate as shown in (20), which is slower than the rate in (18). A broader result than (20), permitting a more general eigen ratio $\lambda_{1}/\lambda_{r}$ for part (ii), is detailed in the appendix. In practice, since the sample covariance tensor includes both the average of signal-by-noise cross-products and the average of noise-by-noise cross-products, it is uncommon to encounter nearly identical sample spiked eigenvalues. Our simulation study demonstrates that while the original composite PCA provides viable initializations when $\lambda_{1}=\lambda_{r}$, its performance is not as good as that of randomized composite PCA using Procedure 2.
@@
In the typical strong factor models where $\lambda_{1}\asymp\lambda_{r}\asymp d$ and $r$ fixed, the rate becomes $O_{\mathbb{P}}(\delta+\sqrt{1/T}+1/d)$, aligning with the convergence rate for the vector factor model when $\delta=0$.
@@
Suppose Assumptions 4.1, 4.2, 4.3 hold. Assume that $\delta_{\max}=\max_{k\leq K}\delta_{k}<1$ with $\delta_{k}$ defined in (16), and $r=O(T)$. Let $1/\gamma=2/\gamma_{1}+1/\gamma_{2}$, $d=d_{1}\cdots d_{K}$, and $d_{\min}=\min_{k\leq K}d_{k}$. Assume $T\leq C\exp\left(d_{\min}^{\vartheta/(2\vartheta+4)}\right)$ and $T\leq C\exp\left((d_{\min}/r)^{\gamma_{1}/2}\right)$. Suppose that for a proper numeric constant $C_{1,K}$ depending on $K$ only, we have

### v2

@@ 3 Estimation
We consider a two-step estimation procedure to derive the loading vectors and latent factors. This approach begins with initialization via RC-PCA, followed by an iterative refinement step utilizing an iterative simultaneous orthogonalization (ISO) procedure.
@@
where $\Theta_{ij}=w_{i}w_{j}{\mathbb{E}}\left[f_{it}f_{jt}\right]$. Its sample analogue, denoted as $\widehat{\Sigma}$, is computed as the average outer product over $T$ observations:
@@
where $\Theta=W(\mathbb{E}F_{t}F_{t}^{\top})W$, $F_{t}=(f_{1t},\cdots,f_{rt})^{\top}$, $W=\operatorname{diag}(w_{1},...,w_{r})$. This unfolding enables classical PCA estimation if the columns of the loading matrix $A$ are orthogonal. Our framework accommodates general non-orthogonal $a_{i}$’s and hence the PCA procedure introduces a bias component, which motivates the second stage refinement. The accuracy of the PCA estimator hinges on the maximum correlation among the loading vectors. When the additional orthogonality condition is imposed as in Babii et al., (2022), the maximum correlation reduces to 0 and hence bias disappears. The first step, termed initialization via RC-PCA, is detailed in Algorithm 1.
To further relax the eigengap assumption imposed in Babii et al., (2022) and Han et al., (2024), we incorporate randomized projection (Procedure 2) into RC-PCA approach [^8] Closely spaced eigenvalues do not imply that the covariance matrix is singular. A covariance matrix can remain full rank even when some of its eigenvalues are identical or nearly identical. In our theoretical analysis, “closely spaced eigenvalues” specifically refers to identical or nearly identical eigenvalues of the population covariance matrix of the noiseless data, defined as $\Theta=W(\mathbb{E}F_{t}F_{t}^{\top})W$, where $F_{t}=(f_{1t},\cdots,f_{rt})^{\top}$, $W=\operatorname{diag}(w_{1},...,w_{r})$. The estimation error for an individual eigenvector depends on the corresponding eigen-gap (Yu et al.,, 2015). When this gap is small, or when the factor number $r$ grows and $\lambda_{1}\asymp\lambda_{r}$, the performance of standard PCA deteriorates, whereas our RC-PCA remains effective. Moreover, the theoretical investigation of RC-PCA under relaxed assumptions is non-trivial and holds independent value.. Random projection, also known as random slicing (Anandkumar et al., 2014b, ; Sun et al.,, 2017) is a well-recognized initialization method in noiseless tensor CP decomposition, which accommodates closely spaced eigenvalues. We extend this approach to the tensor CP factor model. In Algorithm 1 and Procedure 2, the tuning parameters $c_{0}$ and $\nu$ are not highly sensitive to their specific values. In both the simulation and empirical studies presented in the main text, we use the default settings $c_{0}=0.1$ and $\nu=0.8$, and we find that small or moderate deviations from these values do not affect the results [^9] The parameter $c_{0}$ is related to the noise magnitude. Since the randomized projection step guarantees consistency under orthogonal loadings regardless of whether the eigen-gap condition holds, using a larger $c_{0}$ is a safe choice. The parameter $\nu$ is a pre-specified threshold used to filter out tuples that are too similar to the tuple already selected. Each tuple consists of vectors that estimate factor loadings for an unknown rank $i\in\{1,\ldots,r\}$. Therefore, the role of $\nu$ is to ensure that redundant tuples corresponding to the same factor are removed. Further simulation studies regarding $c_{0}$ and $\nu$ can be found in Appendix C.. Note that the condition $\min\{|\widehat{\lambda}_{i}-\widehat{\lambda}_{i-1}|,|\widehat{\lambda}_{i}-\widehat{\lambda}_{i+1}|\}>c_{0}\widehat{\lambda}_{r}$ is checked within Algorithm 1 and automatically determines the appropriate regime based on information from the data.
Input : The observations ${\cal Y}_{t}\in\mathbb{R}^{d_{1}\times\cdots\times d_{K}}$, t=1,,Tt=1,...,T, the number of factors $r$, small constant $0<c_{0}<1$.
@@
[…]
@@
Input : Noisy tensor $\Xi\in\mathbb{R}^{d_{1}\times\cdots\times d_{K}\times d_{1}\times\cdots\times d_{K}}$, rank $s$, number of random projections $L$, mode $h$ for random projection, tuning parameter $\nu$.
@@
2If $h=\emptyset$, pick $h=\arg\max_{1\leq k\leq K}(d_{k})$
3for $\ell=1$ to $L$ do
4   Randomly draw a $d_{h}\times d_{h}$ Gaussian matrix $\theta$ whose entries are i.i.d. $N(0,1)$.
5  Compute $\Xi\times_{h}\times_{K+h}\theta$ and compute its leading singular value and left singular vector $\eta_{\ell},\widetilde{u}_{\ell}$.
6  Compute $\widetilde{a}_{\ell k}$ as the top left singular vector of ${\rm mat}_{k}(\widetilde{u}_{\ell})\in\mathbb{R}^{d_{k}\times(d/(d_{k}d_{h}))}$, for all $k\neq h$.
7  Compute $\widetilde{a}_{\ell h}$ as the top left singular vector of $\Xi\times_{k\neq h,K+h}\widetilde{a}_{\ell k}$, with $\widetilde{a}_{\ell,k}=\widetilde{a}_{\ell,k-K}$ for $k>K$.
@@
Remark 3.1.
To illustrate the idea behind Procedure 2, consider a matrix time series case with $K=2$. If the collinearity or coherence among the CP loading vectors $(a_{1k},...,a_{rk})$ is low, selecting a random projection vector $b$ for the first mode that is close to $a_{11}$ will result in $|a_{i1}^{\top}b|$ being small for all $i>1$ and $|a_{11}^{\top}b|\approx 1$. Consequently, the projected data ${\cal Y}_{t}b=\sum_{i=1}^{r}w_{i}f_{it}(a_{i1}^{\top}b)a_{i2}+{\cal E}_{t}b$ retains the signal corresponding to the first CP factor almost unchanged while significantly reducing the signals from the other factors.
Following initialization, we refine the estimation using an ISO procedure (Algorithm 3). This step aims to enhance estimation accuracy and extract latent factors. The procedure is motivated by the vector factor structure of the denoised ${\cal Y}_{t}$:
${\cal Z}_{t,ik}=w_{i}f_{it}a_{ik}+{\cal V}_{t,ik},$ (8)
@@
$B_{k}=A_{k}(A_{k}^{\top}A_{k})^{-1}=(b_{1k},...,b_{rk})\in\mathbb{R}^{d_{k}\times r}$, $A_{k}=(a_{1k},\ldots,a_{rk})\in\mathbb{R}^{d_{k}\times r}$, and we have used the fact that $b_{ik}$ is orthogonal to all $a_{jk},j\neq i$ by construction. The orthogonalization projection, which takes place in all except the $k$th mode simultaneously in each computational iteration, transforms the tensor ${\cal Y}_{t}$ to a $d_{k}\times 1$ vector, reducing dimensions and noise substantially. This transformation enables easy and accurate estimation of the classical vector factor model in equation (8).
Consider a simple illustrative example with $K=2$, $r=2$:
$\mathcal{Y}_{t}=w_{1}f_{1t}a_{11}\otimes a_{12}+w_{2}f_{2t}a_{21}\otimes a_{22}+\mathcal{E}_{t}=w_{1}f_{1t}a_{11}a_{12}^{\top}+w_{2}f_{2t}a_{21}a_{22}^{\top}+\mathcal{E}_{t}\in\mathbb{R}^{d_{1}\times{d_{2}}}.$
The ideal projection with true $b_{12}$ yields
${\cal Z}_{t,11}$ $={\cal Y}_{t}\times_{2}b_{12}={\cal Y}_{t}\ b_{12}$
$=w_{1}\ f_{1t}\ a_{11}(\underbrace{a_{12}^{\top}b_{12}}_{1})+w_{2}\ f_{2t}\ a_{21}(\underbrace{a_{22}^{\top}b_{12}}_{0})+{\cal E}_{t,11}^{*}$
$=w_{1}\ f_{1t}a_{11}+{\cal E}_{t,11}^{*},$
where ${\cal E}_{t,11}^{*}={\cal E}_{t}\times_{2}b_{12}\in\mathbb{R}^{d_{1}}$. This projection reduces the dimension from $\mathbb{R}^{d_{1}\times{d_{2}}}$ to $\mathbb{R}^{d_{1}}$, while simultaneously increasing the signal-to-noise ratio.
In practice, we do not observe $b_{ik}$ and iterations can be applied to update the estimations. Given the previous estimates $\widehat{a}_{ik}^{(m-1)}$, where $m$ is the iteration number, ${\cal Y}_{t}$ can be denoised via
@@
for t=1,,Tt=1,...,T, and consequently, updated loading vectors $\widehat{a}_{ik}^{(m)}$ are obtained through eigenanalysis based on the contemporary covariance $\widehat{\Sigma}({\cal Z}_{1:T,ik}^{(m)})=\frac{1}{T}\sum_{t=1}^{T}{\cal Z}_{t,ik}^{(m)}{\cal Z}_{t,ik}^{(m)\top}$. The iteration continues until convergence or the maximum number of iterations is reached. The default accuracy is set to $\epsilon=10^{-5}$; a smaller value indicates more precise estimation, which typically results in a larger number of iterations. When our proposed RC-PCA is used as the first-stage (initial) estimator, Algorithm 3 converges very rapidly. In our simulation and empirical applications, we set the maximum number of iterations to $M=100$, but convergence is typically achieved in fewer than 5 iterations.
Input : The observations ${\cal Y}_{t}\in\mathbb{R}^{d_{1}\times\cdots\times d_{K}}$, t=1,,Tt=1,...,T, the number of factors $r$, the warm-start initial estimates $\widehat{a}_{ik}^{(0)}$, $1\leq i\leq r$ and $1\leq k\leq K$, the tolerance parameter $\epsilon>0$, and the maximum number of iterations $M$.
@@ 10        Given previous estimates $\widehat{a}_{ik}^{(m-1)}$, calculate
for t=1,,Tt=1,...,T. Let $\widehat{\Sigma}\left({\cal Z}_{1:T,ik}^{(m)}\right)=\frac{1}{T}\sum_{t=1}^{T}{\cal Z}_{t,ik}^{(m)}{\cal Z}_{t,ik}^{(m)\top}$.
@@
$\widehat{w}_{i}$ $=\left(\frac{1}{T}\sum_{t=1}^{T}\left({\cal Y}_{t}\times_{k=1}^{K}\widehat{b}_{ik}^{(m)\top}\right)^{2}\right)^{1/2},\quad\quad i=1,...,r,$
$\widehat{f}_{it}$ $=\widehat{w}_{i}^{-1}\cdot{\cal Y}_{t}\times_{k=1}^{K}\widehat{b}_{ik}^{(m)\top},\quad i=1,...,r,\ \ t=1,...,T,$
@@
Algorithm 3 Iterative Simultaneous Orthogonalization
@@
Input : The observations ${\cal Y}_{t}\in\mathbb{R}^{d_{1}\times\cdots\times d_{K}}$, t=1,,Tt=1,...,T, the upper bound of the number of factors $r_{\max}$.
@@
[…]
@@
Input : The observations ${\cal Y}_{t}\in\mathbb{R}^{d_{1}\times\cdots\times d_{K}}$, t=1,,Tt=1,...,T, the upper bound of the number of factors $r_{\max}$.
@@ 5  Obtain $\widehat{r}_{k}$ by
Another closely related approach is tensor PCA (TPCA) proposed in Babii et al., (2022). They consider a CP tensor factor model with orthogonal loading vectors. Unlike Tucker factor models, the identification of CP factor models does not necessarily require orthogonality. Applying TPCA to models with non-orthogonal loadings introduces a bias component of higher order than our first-stage RC-PCA. Even when the loadings are orthogonal, our contemporary variance-based iterative estimation exhibits a faster convergence rate than TPCA due to dimension and noise reduction. A comparison of our estimator with the autocovariance-based estimator and TPCA through simulation will be presented in Section 5.
@@ 4 Theory
In this section, we study the statistical properties of the algorithms introduced previously. Our theoretical framework offers guarantees for consistency and outlines the statistical error rates for estimating the factor loading vectors $a_{ik}$, where $1\leq i\leq r,1\leq k\leq K$, given certain regularity conditions. Considering that the loading vector $a_{ik}$ can only be identified with a change in sign, we employ
$\|\widehat{a}_{ik}\widehat{a}_{ik}^{\top}-a_{ik}a_{ik}^{\top}\|_{2}=\sqrt{1-(\widehat{a}_{ik}^{\top}a_{ik})^{2}}=\sup_{\boldsymbol{z}\perp a_{ik}}|\boldsymbol{z}^{\top}\widehat{a}_{ik}|=\|\sin\angle(\widehat{a}_{ik},a_{ik})\|_{2}$
to quantify the discrepancy between $\widehat{a}_{ik}$ and $a_{ik}$. This measure provides a meaningful and practical way to assess the discrepancy between the estimated and true loadings, and we shall apply it in our simulation as well.
@@
Let $F_{t}=(f_{1t},...,f_{rt})^{\top},W=\operatorname{diag}(w_{1},...,w_{r})$, and $\Theta=W(\mathbb{E}F_{t}F_{t}^{\top})W$. Assume that the eigenvalues of $\mathbb{E}(F_{t}F_{t}^{\top})$ satisfy $c_{0}^{-1}\leq\lambda_{r}(\mathbb{E}F_{t}F_{t}^{\top})\leq\cdots\leq\lambda_{1}(\mathbb{E}F_{t}F_{t}^{\top})\leq c_{0}$ for a constant $c_{0}$. Define $\lambda_{i}=\lambda_{i}(\Theta)$ and, without loss generality, let $\lambda_{1}\geq\lambda_{2}\geq\cdots\geq\lambda_{r}>0$. For any $v\in\mathbb{R}^{r}$ with $\|v\|_{2}=1$,
$\mathbb{P}\left(\left|v^{\top}F_{t}\right|\geq x\right)\leq c_{1}\exp\left(-c_{2}x^{\gamma_{1}}\right),$ (14)
@@
Assume the factor process $f_{it},1\leq i\leq r$, is stationary and $\alpha$-mixing in $t$. The mixing coefficient satisfies
@@
Assumption 4.1 allows for cross-sectional dependence in errors and is closely aligned with the noise conditions presented in seminal works such as Bai and Ng, (2002), Bai, (2003), Lam et al., (2011), Lam and Yao, (2012), and others within the factor model literature. It requires that $\xi_{it}$ has exponential-type tails, allowing us to apply large deviation theory. This is a standard condition in the literature on ultrahigh-dimensional data analysis (see, e.g., Chang et al., (2023); Fan et al., (2013)). When $\vartheta=2$, the condition corresponds to a sub-Gaussian tail. Although it might be possible to replace this assumption with finite moment conditions through a significantly more involved theoretical analysis, we have chosen to retain it to focus on the essentials. For simplicity, we assume that the noise tensor remains independent across time $t$, allowing for weak cross-sectional dependence. While incorporating weak temporal correlation among the noise, as suggested by Bai and Ng, (2002), is plausible, it substantially complicates our theoretical analysis. Therefore, we defer this exploration to future research. Nonetheless, our simulation studies show that the proposed methods remain robust even when the noise exhibits weak temporal dependence.
Assumption 4.2 ensures the unique identification of all factor loading vectors $a_{ik}$ up to sign changes. Unlike the eigen decomposition of a matrix, if some $\lambda_{i}$ are equal, the estimation of the loading vectors $a_{ik}$ isn’t subject to rotational ambiguity but only to the signed permutation of loading vectors. Consistent with the classical literature on factor models (e.g., Assumption A in Bai and Ng, (2002), Assumption M(d) in Stock and Watson, (1998)), Assumption 4.2 allows factors to be arbitrarily correlated. Furthermore, Assumption 4.2 specifies that the tail probability of $f_{it}$ must exhibit exponential decay. Specifically, when $\gamma_{1}=2$, it implies that $f_{it}$ follows a sub-Gaussian distribution. We exclude the case of identical factors. In extreme cases, such as when $f_{1t}=f_{2t}$, our composite PCA method (steps 3 and 4 in Algorithm 1) will fail. However, based on Remark 3.1, the random projection method (Procedure 2) should still yield reasonable initializations. We do not pursue this direction further in our theoretical analysis, as it is beyond the scope of our current paper.
@@
Recall $A_{k}$ defined in equation (10) with $a_{ik}$ as its columns. As $\|a_{ik}\|_{2}^{2}=1$, the correlation among columns of $A_{k}$ can be measured by
@@
to measure the correlation of the matrix $A=(a_{1},\ldots,a_{r})\in\mathbb{R}^{d\times r}$ with $a_{i}=\mbox{vec}(\otimes_{k=1}^{K}a_{ik})$ and $d=\prod_{k=1}^{K}d_{k}$. Let $\delta_{\max}=\max\{\delta_{1},\cdots\delta_{K}\}$. Using properties of the Kronecker product, we can show that $\delta\leq\prod_{k=1}^{K}\delta_{k}<\min_{1\leq k\leq K}\delta_{k}\leq\delta_{\max}$.
@@
Suppose Assumptions 4.1, 4.2, 4.3 hold. Let $1/\gamma=2/\gamma_{1}+1/\gamma_{2}$, and $\delta<1$ with $\delta$ defined in (17).
(i). The eigengaps satisfy $\min\{\lambda_{i}-\lambda_{i+1},\lambda_{i-1}-\lambda_{i}\}\geq c\lambda_{r}$ for all $1\leq i\leq r$, with $\lambda_{0}=\infty,\lambda_{r+1}=0$, and $c$ is sufficiently small constant. With probability at least $1-T^{-C_{1}}-d^{-C_{1}}$, the following error bound holds for the estimation of the loading vectors $a_{ik}$ using Algorithm 1,
@@
$\phi^{(0)}$ $=\lambda_{1}\left(\sqrt{\frac{r+\log T}{T}}+\frac{(r+\log T)^{1/\gamma}}{T}\right)+\sqrt{\frac{\lambda_{1}d\log d}{T}}+\frac{\sqrt{\lambda_{1}dr}\log(d)(\log T)^{1+\frac{2}{\vartheta}+\frac{1}{\gamma_{1}}}+d\log(d)(\log T)^{\frac{2\vartheta+4}{\vartheta}}}{T}+1.$ (19)
@@
The first term of the upper limit in (18) and (20) arises from the non-orthogonality of the loading vectors $a_{ik}$, which can be interpreted as bias. Meanwhile, the second term in (18) stems from a concentration inequality for random noise and thus reflects a form of stochastic error.
When the eigengap condition is not met, we employ randomized projection to determine the statistical convergence rate as shown in (20), which is slower than the rate in (18). A broader result than (20), permitting a more general eigen ratio $\lambda_{1}/\lambda_{r}$ for part (ii), is detailed in the appendix. In practice, since the sample covariance tensor includes both the average of signal-by-noise cross-products and the average of noise-by-noise cross-products, it is uncommon to encounter nearly identical sample spiked eigenvalues. Our simulation study demonstrates that while the original composite PCA provides viable initializations when $\lambda_{1}=\lambda_{r}$, its performance is not as good as that of RC-PCA using Procedure 2.
@@
In the typical strong factor models where $\lambda_{1}\asymp\lambda_{r}\asymp d$ (i.e. $w_{i}\asymp\sqrt{d}$) and $r$ fixed, the rate becomes $O_{\mathbb{P}}(\delta+\sqrt{1/T}+1/d)$, aligning with the convergence rate for the vector factor model when $\delta=0$.
Remark 4.2.
Procedure 2 provides a theoretically guaranteed strategy for resolving the repeated eigenvalue issue and can, intuitively, also be applied uniformly in Algorithm 1. However, due to the random projection step, the enhanced incoherence parameter $\delta$ is compromised, resulting in slower convergence rates as indicated in (20) compared to (18). In contrast, Auddy and Yuan, (2023) assumes an orthogonal decomposable tensor, thereby avoiding bias issues; for their purpose, they only provide preliminary results with a convergence rate slower than $1/4$ under the same error metric. Similarly, Anandkumar et al., 2014b considers the random projection method, but their analysis is limited to tensor CP decomposition with the eigen-ratio assumed to be of constant order.
Moreover, when the eigen-ratio is large, the number of required initializations increases exponentially with the eigen-ratio (see Theorem 4.1(ii)). From a computational cost perspective, the composite PCA remains valuable.
@@
Suppose Assumptions 4.1, 4.2, 4.3 hold. Assume that $\delta_{\max}=\max_{k\leq K}\delta_{k}<1$ with $\delta_{k}$ defined in (16), and $r=O(T)$. Let $1/\gamma=2/\gamma_{1}+1/\gamma_{2}$. Assume